Solve it
(D^2+DD'-6D'^2)z= ysinx
1
Expert's answer
2021-08-25T11:39:07-0400

Auxiliary equation will be


"m^2+m-6=0"

"(m-2)(m+3)=0"

"m_1=2, m_2=-3"

Hence


"C.F.=\phi_1(y+2x)+\phi_2(y-3x)"

"P.I.=\dfrac{1}{D^2+DD'-6D'^2}y\sin x"

"=I.P.\ of\ e^{ix}\dfrac{1}{(D+i)^2+(D+i)(D'+0)-6(D'-0)^2}y"

"=I.P.\ of\ e^{ix}\dfrac{1}{D^2-1+2iD+DD'+iD'-6D'^2}y"

"=I.P.\ of\ e^{ix}\dfrac{1}{D^2+DD'-6D'^2+i(2D+D')-1}y"

"=I.P.\ of\ -e^{ix}\dfrac{1}{1-[D^2+DD'-6D'^2+i(2D+D')]}y"

Using binomial expansion


"P.I.=I.P.\ of\ -e^{ix}(1+[i(2D+D')+D^2+DD'-6D'^2])y"

"=I.P.\ of\ -e^{ix}(y+i)"

"=I.P.\ of\ -(\cos x-i\sin x)(y+i)"

"P.I.=-\cos x-y\sin x"

Hence the general solution is given by


"y=\phi_1(y+2x)+\phi_2(y-3x)-(\cos x+y\sin x)"